Cash Holdings |87 Calculate JWs CCE, Cash Ratio, DCH, and NLB, and interpret the 2-year th sures. How would you describe JWs liquidity position? Suppose that the JWs treasurer has targeted a DCH of 250 days for the next fiscal year. Assuming that CGS are expected to remain unchanged from 2016, then what (increase or decrease) in cash would be warranted change Assuming that cash holdi or decrease) in CGS would be warranted? b. ngs will remain unchanged from 2016, then what change (increase 3. In an interview on CNBC, d major shareholder of JWs common stock (a known activist inves- tor) stated that JWs management is holding too much cash and that some of this liquidity should be reinvested in the firm. Consequently, the CFO will lower the Cash Ratio by using cash and equivalents to purchase operating assets (i.e., the firms total asset base will not change). What level of cash and equivalents is required to generate the following values for the Cash Ratio? a. b. 20% 15% 10% 5% d. JWs shareholders are once again causing problems. This time, the issue concerns the CCE. Shareholders want to see the firm become more efficient at turning revenues into operating cash flow. Assuming that revenues remain unchanged from fiscal year 2016, what level of operating cash flow will be required to generate the following values for the CCE? a. 10%? b. 15% С. 20% d. 25% ws treasurer has just decided to incorporate the λ concept when evaluating the firms liquid- ity position. In addition to the information provided by the financial statements, what other variables would also be needed to calculate JWs λ?
PRUBLL Use the following financial statements for problems t 2015 5,050.00 4,040.00) $ 1,010.00 $ 420.00) (S 26.00) 394.00 (S 30.00) $ 364.00 (S 145.60) S 218.40 JW, Inc.: Income Statement 2016 5,70000 S 4,560.00) $ 1,140.00 (S 500.00) Sales CGS Gross Profit Operating Expenses Depreciation EBIT Interest Expense EBT Taxes NI S 27.00) $ 613.00 S 35.00) $ 57800 $ 231.20) 346.80 Dividends 5 218.40 $346.80 JW, Inc.: Balance Sheet 2016 Cash and Equivalents Accounts Receivable Inventory Total Current Assets Fixed Assets Accumulated Depreciation Net Fixed Assets $ 50000 $300.00 $120.00 $ 500.00 $ 340.00 $ 1.51 $ 660.00 800.00 $ 433.00 $1,947 80 Total Assets $ 600.00 $ 1,560.00 Accounts Payable Short-Term Notes Payable Accruals Total Current Liabilities Long-Term Debt $11.00 $371.00 780.00 $1,151.00 $ 360.00 Common Stock Retained Earnings Total Shareholders Equity $ 120.00 $ 676.80 Liabilities and Shareholders Equity $1,94780
Answer #1

1 CCE 2016 2015

714.8 120

Trend CCE increased in 2016 by 594.8 as compare to 2015

Cash ratio In 2015 120/360=0.33

Cash ratio In 2016 714.8/371=1.92

Trend-Increasing cash ratio in 2016 as compare to 2015

DCH calculation

Per day operating expenses in 2015

420/365=1.15

CCE in 2015 =120, so 120/1.15=104.34

Per day operating expense in 2016

500/365=1.36

CCE in 2016 =714.8 so 714.8/1.36=525.58

NLB

2015 120-300=180

2016 714.8-320=394.8

2)

If DCH 250 days then case holding will be increased as compare to 2016

If Cash holding unchanged then CGS will be decreased as compare to 2016

3)

a